OpenStudy (anonymous):

The table shows the different packages of toothpaste available at a store. Bert buys cartons of Orange Cheer and Apple Smile toothpaste. What is the least number of cartons of Orange Cheer and Apple Smile that Bert could buy so that he will have the same number of each flavor of toothpaste? items cartons in package orange cheer 8 great grape 10 apple smile 6

OpenStudy (anonymous):

i think it is 24

OpenStudy (anonymous):

PLEASE HELP!!!

OpenStudy (anonymous):

this is a least common multiple problem. The easiest way to solve these is to break each number into it's prime factors then create a new number using the hightest power of the primes. orange cheer \[8=2^{3}\] great grape \[10=2*5\] apple smile \[6=2*3\] so the smallest number that has all of these primes is \[2^{3}*3*5=120\] so we will have 120 of each flavor to find how many packages we need, we divide. orange cheer \[120/8=15\] great grape \[120/10=12\] apple smile \[120/6=20\]

OpenStudy (anonymous):

so 48?

OpenStudy (anonymous):

one would need to buy 15 boxes of orange cheer, 12 boxes of great grape, and 20 boxes of apple smile for a total of 47 boxes.

OpenStudy (anonymous):

these are my choice's 480 48 24 16